Da Silva and Ehlers decided to admit Fischer as from 1 January as a partner under the following terms:

Ficsher will receive 1/5 of the profits and losses.

Da Silva and Ehlers will contribute in the ratio of 1:2 toward Fischers 1/5 profit share.

Fischer must pay N$1800000 for her 1/5 share in the partnership assets.

The assets of the partnership were re -valued on 1 January 2018 and the revaluation account revealed a surplus of N$60000.


Required:

Calculate the new profit sharing ratio after the admission of Fichser.


1
Expert's answer
2022-03-28T14:16:34-0400

Old Profit sharing ratio between Da Silva and Ehlers = 3:2

New partner Fischer will receive "\frac{1}{5}" of profits and losses

Da Silva and Ehlers sacrificing ratio=2:1


From Da Silva, Fischer will take:

"\frac{2}{3} \times \frac{1}{5} = \frac{2}{15}"


From Ehlers, Fischer will take:

"\frac{1}{3} \times \frac{1}{5} = \frac{1}{15}"


Da Silva's new share:

= "\frac{3}{5} - \frac{2}{15}"

="\frac{7}{15}"


Ehlers' new share:

="\frac{2}{5} - \frac{1}{15}"

="\frac{5}{15}"


New profit ratio for Da Silva, Ehlers and Fischer:

="\frac{7}{15}: \frac{5}{15}:\frac{3}{15}"


=7:5:3